OpenStudy (anonymous):

1 2 3 2 1 3 2 -2 0 1 0 1

OpenStudy (farmdawgnation):

A matrix?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Let L : R3 -> R4 be a map defined as follows: if x in R3 then L(x) = Ax. Show that L is a linear map.

OpenStudy (anonymous):

@zarkon

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

So we need to show \[L(x_1+x_2)=A(x_1+x_2)=Ax_1+Ax_2=L(x_1)+L(x_2)\] and \[L(c x_1)=A(c x_1)=c A(x_1)=c L(x_1)\] where c is a number and \[x_1,x_2 \in \mathbb{R}^3\]

OpenStudy (anonymous):

ok

myininaya (myininaya):

\[\text{Let } x_1=(a_1 , a_2, a_3)^t \text{ and let } x_2=(b_1,b_2,b_3)^t \]

myininaya (myininaya):

So ...I don't know how to do matrices in latex (because I suck) But I can draw it

OpenStudy (anonymous):

that's ok , just type it out like I did

myininaya (myininaya):